an object of mass 4 kg is placed on an inclined plane at an angle of 60. if the object slides down the plane with an acceleration of 3m/s, what is the coefficient of the kinetic friction
1
Expert's answer
2019-01-04T10:56:11-0500

The Newton’s second law gives

ma=mg sin⁡θ-F_frict

F_frcit=μN=μmg sin⁡θ

So

a=g sin⁡θ-μg cos⁡θ

μ=(g sin⁡θ-a)/(g cos⁡θ )=(9.8×sin⁡〖60°〗-3)/(9.8 cos⁡〖60°〗 )=1.12